OpenStudy (anonymous):

Help me PLEASE I'm confused!!! The hypotenuse of a right triangle is 1cm longer than the longer leg. The shorter leg is 7cm shorter than the longer leg. Find the length of the longer leg of the right triangle.Label each side of the triangle. Then substitute into Pythagorean Theorem.

hartnn (hartnn):

at which step are u confused ?

OpenStudy (anonymous):

Shorter leg = x longer leg = x+7 hypotenuse = x+7+1 = x+8 Do you know how to proceed now ?

OpenStudy (anonymous):

no I don't know-how to continue

OpenStudy (anonymous):

hartnn i dont know how to figure out what belongs to what leg

hartnn (hartnn):

did u get @Aditi_Singh 's explanation ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I'm guessing that x is the length of the longer leg, and that this can be written as: \[X^2+ (X-7)^2 = X+1 \] Then, you can solve from there by expanding brackets and you get a quadratic equation: \[x^2 -8x+48=0\]

OpenStudy (anonymous):

Almost Correct , but x^2 + (x+7)^2 = (x+8)^2 Apply Pythagoras theorem!

OpenStudy (anonymous):

Wait no.. my answer makes sense. Doesnt it? \[x^2+(x−7)^2=(x+1)^2 \] \[= x^2+x^2-7x+49=x^2 + x+1 \] \[x^2-8x+48=0 \]

OpenStudy (anonymous):

:| , ofcourse it does makes sense. But here, your hypotenuse isn't (x+1) , it is x+8 ..

OpenStudy (anonymous):

Im taking x as the longer leg, not the shorter one

OpenStudy (anonymous):

Okay, then the longer leg would be x , shorter length would be x-7, and hyp would be x+1 .. Well done buddy :D , you were absolutely correct :D

OpenStudy (anonymous):

Woohoo ;) High five for being right :D

OpenStudy (anonymous):

High five ;) :D